# Stony Brook, NY 11794, USA

### v1

@@ 4 $A$-twisting and TQFT modular data
$\mathcal{F}(Z):=e^{[W(Z)-Z_{i}\,\partial_{Z_{i}}\!W(Z)]/(2\pi i)},$ (142)
in terms of which we have the map (see e.g. Gang:2023rei ; Cho:2020ljj ; Gang:2021hrd ) [^10] Instead of identifying $T^{2}$ with $\mathcal{F}^{-2}$, we could identify $T^{2}$ with $\mathcal{F}^{2}$ but either $i)$ consider the conjugate theory corresponding to $\gamma=-1$, or $ii)$ reverse parity by negating our $k_{gg}$ and $k_{gR}$. The result would again match the modular data of $\widehat{\mathfrak{su}}(2)_{-4/3}$, up to the phase/sign ambiguities. Alternatively, we can identify $T^{2}$ with $\mathcal{F}^{2}$ with our current data, but match with the MTC conjugate to that of (the admissible modules of) $\widehat{\mathfrak{su}}(2)_{-4/3}$.
$\{S^{2}_{0\alpha},T^{2}_{\alpha\alpha}\}\longleftrightarrow\{\mathcal{H}(Z^{\ast}_{\alpha})^{-1},\mathcal{F}(Z^{\ast}_{\alpha})^{-2}\}\,.$ (143)
This gives in the limit $a\to 1$ that
@@
with the asterisks undetermined by (143). This partial determination is compatible with (134).
@@
we can fill in the asterisks in (145), compatibly with (134).
@@ 5 Discussion and future directions
Besides the simplest new case of $(A_{1},A_{3}),$ it would be interesting to study the TQFTs arising from the A-twist of the next-to-simplest 3d SCFTs on the second sheet of $(A_{1},D_{3})$ and $(A_{1},D_{4}).$ The case of $(A_{1},D_{3})$ would presumably be similar to that of $(A_{1},A_{3})$, since they are dual to each other, but the case of $(A_{1},D_{4})$ is expected to be quite different, and related to the VOA $\widehat{\mathfrak{su}}(3)_{-3/2}$ Buican:2015ina ; Cordova:2015nma ; Beem:2017ooy . Our preliminary analysis indicates that the corresponding A-twisted TQFT has only a single Bethe root. Identifying the (four Creutzig:2021seo ) bulk line operators associated with the VOA admissible modules seems therefore more challenging for $(A_{1},D_{4})$ than in the $(A_{1},A_{3})$ case.
@@
For the $(A_{1},A_{3})$ EFT, one can use the operator $\epsilon_{ab}\Phi^{a}_{-1}\Phi^{b}_{-1}\Phi_{+2}$ for the superpotential deformation. The resulting IR TQFT is expected to have the same $Z_{g,p}$ partition function (possibly modulo a phase factor) as the $\widehat{\mathfrak{su}}(2)_{-4/3}$ theory, which can be computed using the modular matrices given in (134). Interestingly, $\widehat{\mathfrak{su}}(3)_{-1}$ fulfills this role and can be identified as the IR TQFT. The modular matrices of the $\widehat{\mathfrak{su}}(3)_{-1}$ are
$\begin{split}\tilde{T}&=\begin{pmatrix}e^{\frac{\pi i}{12}}&0&0\\ 0&e^{-\frac{7\pi i}{12}}&0\\ 0&0&e^{-\frac{7\pi i}{12}}\end{pmatrix},\qquad\tilde{S}=\frac{1}{\sqrt{3}}\begin{pmatrix}1&1&1\\ 1&\omega&\omega^{2}\\ 1&\omega^{2}&\omega\end{pmatrix}.\end{split}$ (152)
In this way, we can relate unitary semisimple TQFTs to non-unitary logarithmic TQFTs that share the same $Z_{g,p}$ partition functions, provided that the latter can be realized as A-twists of 3d $\mathcal{N}=4$ SCFTs with a trivial Coulomb branch but a non-trivial Higgs branch. This relationship may offer a simple criterion for determining whether a given non-unitary logarithmic TQFT can be realized as an A-twisted theory of such 3d $\mathcal{N}=4$ SCFTs. It would also be interesting to investigate further this systematic way of constructing unitary TQFTs in 3d $\mathcal{N}=4$ SCFTs with trivial Coulomb branch but non-trivial Higgs branch, for example relating the specific resulting TQFT to the geometric properties of the Higgs branch only.
@@ Acknowledgements.
where the quotient acts as the center of the $SU(2)$ factor and the transformation $\mathrm{e}^{i\pi}$ of the $U(1)$. The matter content is

### v2

@@ 4 $A$-twisting and TQFT modular data
$\frac{(a+z_{\ast})(az_{\ast}+1)}{a\left(z_{\ast}^{2}-1\right)^{2}}=1.$ (139)
We discard the root $z_{\ast}=0$ since it yields a divergent handle-gluing operator; see Eq. (142) below. The Bethe equation can then be simplified to
@@
$\mathcal{F}(Z):=e^{[W(Z)-Z_{i}\,\partial_{Z_{i}}\!W(Z)]/(2\pi i)}.$ (143)
In terms of these we have the map (see e.g. Gang:2023rei ; Cho:2020ljj ; Gang:2021hrd ) [^11] Instead of identifying $T^{2}$ with $\mathcal{F}^{-2}$, we could identify $T^{2}$ with $\mathcal{F}^{2}$ but either $i)$ consider the conjugate theory corresponding to $\gamma=-1$, or $ii)$ reverse parity by negating our $k_{gg}$ and $k_{gR}$. The result would again match the modular data of $\widehat{\mathfrak{su}}(2)_{-4/3}$, up to the phase/sign ambiguities. Alternatively, we can identify $T^{2}$ with $\mathcal{F}^{2}$ with our current data, but match with the MTC conjugate to that of $\widehat{\mathfrak{su}}(2)_{-4/3}$.
$S^{2}_{0\alpha}\longleftrightarrow\mathcal{H}(Z^{\ast}_{\alpha})^{-1},\qquad T^{2}_{\alpha\alpha}\longleftrightarrow\mathcal{F}(Z^{\ast}_{\alpha})^{-2}\,,$ (144)
up to $\alpha$-independent overall phases. This gives in the limit $a\to 1$ that
@@
[…]
@@
we fill in the asterisks in (146), compatibly with (134), modulo the phase ambiguities.
@@ 5 Discussion and future directions
Besides the simplest new case of $(A_{1},A_{3}),$ it would be interesting to study the TQFTs arising from the A-twist of the next-to-simplest 3d SCFTs on the second sheet of $(A_{1},D_{3})$ and $(A_{1},D_{4}).$ The case of $(A_{1},D_{3})$ should be similar to that of $(A_{1},A_{3})$, since they are dual to each other, but the case of $(A_{1},D_{4})$ is expected to be quite different, and related to the VOA $\widehat{\mathfrak{su}}(3)_{-3/2}$ Buican:2015ina ; Cordova:2015nma ; Beem:2017ooy . The 3d $\mathcal{N}=2$ data appropriate for the A-twist in the latter case can be found by setting $s=1$ in Table 8. The result is displayed in Table 11.
$U(1)_{g}$ $SU(2)_{f}$ $U(1)_{v}$ $U(1)_{R_{s=1}}$
$\Phi_{1}$ $1$ $\bf 1$ 1 1
$\tilde{\Phi}_{1}$ $1$ $\bf 1$ $-1$ 1
$\Phi_{-1}$ $-1$ $\bf 2$ 0 1
$\Phi_{2}$ $2$ $\bf 1$ 0 0
$\Phi_{-2}$ $-2$ $\bf 1$ 0 0
$\beta$ 0 $\bf 1$ 0 $2$
Table 11: The 3d $\mathcal{N}=2$ data of the A-twist $\text{TQFT}^{\,\gamma=1}_{(A_{1},D_{4})}$. All CS levels vanish.
Our preliminary analysis indicates that the corresponding TQFT has only a single Bethe root. Identifying the (four Creutzig:2021seo ) bulk line operators associated with the VOA admissible modules seems therefore more challenging for $(A_{1},D_{4})$ than in the $(A_{1},A_{3})$ case.
@@
For the $(A_{1},A_{3})$ EFT, one can use the operator $\epsilon_{ab}\Phi^{a}_{-1}\Phi^{b}_{-1}\Phi_{+2}$ for the superpotential deformation. The resulting IR TQFT is expected to have the same $Z_{g,p}$ partition function (possibly modulo a phase factor) as the $\widehat{\mathfrak{su}}(2)_{-4/3}$ theory, which can be computed using the modular matrices given in (134). Interestingly, the conjugate (or the time-reversed) TQFT of $\widehat{\mathfrak{su}}(3)_{1}$ fulfills this role and can be identified as the IR TQFT. The modular matrices of $\widehat{\mathfrak{su}}(3)_{1}$ are
$\begin{split}{T}_{\widehat{\mathfrak{su}}(3)_{1}}&=\begin{pmatrix}e^{-\frac{\pi i}{6}}&0&0\\ 0&e^{\frac{\pi i}{2}}&0\\ 0&0&e^{\frac{\pi i}{2}}\end{pmatrix},\qquad{S}_{\widehat{\mathfrak{su}}(3)_{1}}=\frac{1}{\sqrt{3}}\begin{pmatrix}1&1&1\\ 1&\omega&\omega^{2}\\ 1&\omega^{2}&\omega\end{pmatrix}.\end{split}$ (153)
${T}_{\widehat{\mathfrak{su}}(3)_{1}}$ and $T_{\widehat{\mathfrak{su}}(2)_{-4/3}}$ are indeed complex conjugate modulo an overall phase, while ${S}_{\widehat{\mathfrak{su}}(3)_{1}}$ and $S_{\widehat{\mathfrak{su}}(2)_{-4/3}}$ differ by complex conjugation up to an overall sign, as well as negating the simple object corresponding to the second row/column. [^13] In the rational context of ArabiArdehali:2024ysy , imposing positivity of the fusion coefficients could limit the sign flip possibilities of the simple objects in that setting. The overall phase and sign differences, and the sign flip of the simple object, are all accommodated by the unresolved ambiguities in (144), and manifest themselves as overall phase differences in the partition functions (152).
In this way, we can relate unitary semisimple TQFTs to non-unitary logarithmic TQFTs that share the same $Z_{g,p}$ partition functions (up to overall phases), provided that the latter can be realized as A-twists of 3d $\mathcal{N}=4$ SCFTs with a trivial Coulomb branch but a non-trivial Higgs branch. This relationship may offer a simple criterion for determining whether a given non-unitary logarithmic TQFT can be realized as an A-twisted theory of such 3d $\mathcal{N}=4$ SCFTs. It would also be interesting to investigate further this systematic way of constructing unitary TQFTs in 3d $\mathcal{N}=4$ SCFTs with trivial Coulomb branch but non-trivial Higgs branch, for example relating specific observables in the TQFTs to the geometric properties of the Higgs branch only.
@@ Acknowledgements.
where the quotient acts as the center of the $SU(2)$ factor and the transformation $\mathrm{e}^{i\pi}$ of the $U(1)$. The matter content is displayed in Table 13.